OpenStudy (anonymous):

Simplify the expression where possible. (r ^3) ^-2

OpenStudy (shamil98):

(r^3)^-2 = \[\frac{ 1 }{ (r^3)^2 }= \frac{ 1 }{ r^6 }\]